# Empathy Level Alignment via Reinforcement Learning for Empathetic Response Generation

### v2

@@ IV-D Evaluation Metrics
Human Rating: We randomly sample 100 dialogue contexts and their corresponding responses generated by EmpRL and baseline models. Three human evaluators are then asked to rate the responses in terms of three aspects:
@@
Ratings for each aspect are on a scale from 1 to 5, with higher scores indicating better performance (i.e., 1: poor, 2: marginal, 3: moderate, 4: good, 5: excellent). The average score from the three evaluators is used as the final rating.
Human A/B Test: To compare our EmpRL framework with other baselines, we also conduct human A/B tests. Model A represents EmpRL, while model B corresponds to a baseline model. In the test, three evaluators are asked to choose which model they believe performs better between A and B. If they find both models equally effective, they can choose “Tie”. The final result is determined by the majority vote. In cases where the three evaluators provide conflicting results, a fourth evaluator is introduced. The proportions of “Win”, “Lose”, and “Tie” for model A are reported.
@@ V-A Automatic Evaluation Results
The automatic evaluation results on EmpatheticDialogues are presented in Table II. From the table, we conclude that: (1) Models such as CASE, SEEK, and CEM, which incorporate both affective and cognitive empathy, outperform models that focus solely on affective empathy, such as MoEL, MIME, and KEMP. This validates that considering both types of empathy leads to better response generation. (2) In terms of PPL, EmpRL shows a reduction of $22.79$ compared to the state-of-the-art baseline CASE, indicating that leveraging pre-trained language models can generate more fluent responses. Llama3 achieves a lower PPL primarily because it is an LLM with more extensive training datasets and parameters. (3) Regarding Dist-1/2, EmpRL performs comparably to Llama3 and ChatGPT, demonstrating that the proposed EmpRL is capable of generating diverse responses. (4) In terms of Emp-F1, EmpRL achieves the highest score of $69.58\%$. This result indicates that its responses exhibit a greater similarity in empathy levels to the target responses, further validating the effectiveness of EmpRL.
TABLE II: Automatic evaluation results on EmpatheticDialogues.
Models PPL Dist-1 Dist-2 Emp-F1
@@
TABLE III: Automatic evaluation results on PEC.
Models PPL Dist-1 Dist-2 Emp-F1
@@ V-B Human Evaluation Results
TABLE IV: Human rating results on EmpatheticDialogues.
@@
[…]
On the other hand, through manually inspecting the responses generated by ChatGPT, we reach a conclusion similar to Zhao et al.[52]: ChatGPT frequently repeats the pattern of “emotional restatement + information expansion” when expressing empathy. This repetitive pattern may cause users to experience monotony and a lack of freshness. Therefore, enhancing personalized empathy capabilities requires greater attention.
TABLE V: Human A/B test results on EmpatheticDialogues.

### v3

@@ IV-D Evaluation Metrics
Human Evaluation: We randomly sample 100 dialogue contexts along with their corresponding responses generated by EmpRL and baseline models. Three graduate students with expertise in natural language processing and affective computing from our research team are recruited as human evaluators. To ensure consistency and reliability, all evaluators receive detailed instructions and complete a brief practice session before assessments. This preparatory phase familiarizes them with the evaluation process and align their interpretations of the evaluation criteria. Human evaluations are conducted under a double-blind setup to mitigate potential biases. The specific instructions for both the human rating and human A/B test are outlined below.
Human Rating: The quality of generated responses is assessed based on three criteria: empathy, relevance, and fluency. Ratings for each criterion are on a scale from 1 to 5, where higher scores indicate better performance (1: poor, 2: marginal, 3: moderate, 4: good, 5: excellent).
@@
The average score from three evaluators is used as the final rating.
Human A/B Test: Given a dialogue context and two candidate responses (A and B), please select the preferred response through a comprehensive assessment of their overall quality. This assessment should take into account multiple dimensions, including empathetic expression, contextual relevance, and linguistic fluency. If both responses are deemed equally effective, a “Tie” option is available.
The final result is determined by majority vote. In cases where the three evaluators provide conflicting results, a fourth evaluator is introduced. The proportions of “Win”, “Lose”, and “Tie” outcomes for EmpRL, in comparison to the baselines, are reported.
@@ V-A Automatic Evaluation Results
The automatic evaluation results on EmpatheticDialogues are presented in Table II. We conduct EmpRL with $5$ random seeds and report the mean scores along with $95\%$ confidence intervals. From the table, we conclude that: (1) Models such as CASE, SEEK, and CEM, which incorporate both affective and cognitive empathy, outperform models that focus solely on affective empathy, such as MoEL, MIME, and KEMP. This validates that considering both types of empathy leads to better response generation. (2) In terms of PPL, EmpRL shows a reduction of $22.69$ compared to the state-of-the-art baseline CASE, indicating that leveraging pre-trained language models can generate more fluent responses. Llama3 achieves a lower PPL primarily because it is an LLM with more extensive training datasets and parameters. (3) Regarding Dist-1/2, EmpRL performs comparably to Llama3 and ChatGPT, demonstrating that the proposed EmpRL is capable of generating diverse responses. (4) In terms of Emp-F1, EmpRL achieves the highest score of $69.43\%$. This result indicates that its responses exhibit a greater similarity in empathy levels to the target responses, further validating the effectiveness of EmpRL.
TABLE II: Automatic evaluation results on EmpatheticDialogues, with confidence intervals reported. $\uparrow$ means a higher score is better whereas $\downarrow$ is exactly the opposite. Numbers in bold denote the best results. $\dagger$, $\ddagger$ represent that the improvement over the best and second-best baselines is statistically significant (one-sample t-test, $p$-value $<0.05$), respectively.
[…]
@@
TABLE III: Automatic evaluation results on PEC, with confidence intervals reported. $\dagger$, $\ddagger$ represent that the improvement over the best and second-best baselines is statistically significant (one-sample t-test, $p$-value $<0.05$), respectively.
[…]
@@ V-B Human Evaluation Results
TABLE IV: Human rating results on EmpatheticDialogues. $\ddagger$ represents that the improvement over the second-best baseline is statistically significant (independent t-test, $p$-value $<0.01$), and $95\%$ confidence intervals are reported.
@@
$\pm$0.10 $\pm$0.10 $\pm$0.07
TABLE V: Human A/B test results on EmpatheticDialogues. $\dagger$, $\ddagger$ represent significant improvement with $p$-value $<0.01/0.05$ (sign test), respectively.
@@
[…]
On the other hand, as observed by Zhao et al.[52], ChatGPT often repeats the pattern of “emotional restatement + information expansion” when expressing empathy. To further validate this observation, we perform a manual inspection of the responses generated by ChatGPT and EmpRL for human evaluation samples, assessing whether their responses exhibit this repetitive pattern. The results, including the repetition proportion (defined as the percentage of responses containing repetitive patterns) and the non-repetition proportion, are presented in Fig. 6. We find that the proportion of ChatGPT repeating this pattern is as high as $47.0\%$. This repetitive pattern may cause users to experience monotony and a lack of freshness. Therefore, enhancing personalized empathy capabilities requires greater attention.
(a) ChatGPT
(b) EmpRL
Figure 6: Proportions of repetitive and non-repetitive patterns across models on EmpatheticDialogues.